RC meeting and Run Plan, Friday Oct. 12

From HallCWiki
Jump to: navigation, search


  • Beam quality
    • Bottom line: things were mostly fine during swing until 11pm
    • Current fluctuating at 40uA until we moved SHMS, then we requested 35uA\
    • BLM trip investigation for 6 hours
    • Just started data taking at 35uA again
  • Progress
    • Finished kinematic point (theta_Shms=9.42) after 48+ hours with 46%, short of 50-60% goal
    • Moved to new kinematic point (theta_Shms=6.495), need to limit current to 35uA

Issues: Resolved

Issues: Requiring attention


  • SHMS at small angle (part 1)
    • SHMS raw singles rates very high
      • Found to be around 700kHz at 35uA --> 80% tracking efficiency
      • Briefly ran at 30uA to see increase in efficiency when running at 15% less current --> 83% tracking efficiency (~600kHz singles rate)
      • Decided to run at 35uA
      • Took brief run during beam-tuning at 40uA --> 78% tracking efficiency (820kHz singles rate)
  • SHMS DC1 tripping during beam studies
    • SHMS DC1 tripped when beam studies reached 50uA
    • Decided to leave HV off until we are back to production running
    • Instructed swing shift leader to switch back on if they (finally) finish their beam studies
    • This limits what we can do during the luminosity scan
  • Background rates higher than initialy expected (Garth)


  • SHMS at small angle (part 2)
    • Added during RC meeting and Run Plan, Thursday Oct. 11
    • (Might) need new limits on diffusor ion chamber if we run above >=50uA (Jay Benesh confirmed might be the case), more info here.
      • Not an issue right now because we are rate-limited to 35uA,
      • Might become issue at the next momentum setting and/or the luminosity scan

Request (non-urgent)

  • RadCon wants to put a small detector in the SHMS hut during the next opportunity
    • Needs few hours, communicate with Brad when next opportunity arrises
  • Hall D needs opportunistic 1shift of HallC down for beam studies, not pressing right now (needs to be done before end-of-run)
    • Unlikely this will happen in the next week

Plans - Physics

  • Continue for ~5 more shifts with Q2=3. W=3.14 x=0.25 theta_Shms=6.495
    • might only reach 20-25% of statistics goal
  • Next setting (during weekend): Q2=5.5 W=3.02 x=0.4 theta_Shms=6.495
  • Prepare for 1-2 shifts of luminosity scan early next week
    • Proposed to Jay to do Monday and/or Tuesday SWING when we have experienced shift crew
    • Might be prudent to recruit third worker